OpenStudy (anonymous):

help please! trig identities! use algebra to solve each equation over the domain of -2pi < x < 2 pi a) 2sec^2x - 2 = 3 + tan x b) 4sinxcosx=2 - 2cos^2x

OpenStudy (tkhunny):

So, you're saying you don't know a Pythagorean identity using secant and tangent?

OpenStudy (anonymous):

i do it sec^2 + tan^2 = 1

OpenStudy (anonymous):

you replace it

OpenStudy (tkhunny):

No, that's still incorrect, just like the last time. \(\sin^{2}(x) + \cos^{2}(x) = 1\) \(\tan^{2}(x) + 1 = \sec^{2}(x)\)

OpenStudy (anonymous):

ok

OpenStudy (tkhunny):

Use that to get rid of the secant.

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

ok what about b)?

OpenStudy (tkhunny):

How are your double angle formulas? Are they in your head?

OpenStudy (anonymous):

i have a formula sheet

OpenStudy (anonymous):

it's provided on the test as well

OpenStudy (tkhunny):

Divide the whole things by 2 then take a good, hard look at both sine and cosine double angle formulas.

OpenStudy (anonymous):

can u work out a) for me please. i'm still not getting the right answers.. i thought it was tan x =3/2 and -2.. in radians or degrees but that doesn't seem like it

OpenStudy (tkhunny):

It would help if you would SHOW what you are doing. 4sinxcosx=2 - 2cos^2x I suggested that you divide by 2. Did you do that?

OpenStudy (anonymous):

a) 2x-3=0 x+2=0 b) how do u do sin2x=sin^2x

OpenStudy (anonymous):

for a) i got the answers of .98, and 4.12 but couldn't get the -2 answer